Preble Company manufactures one product. Its variable manufacturing overhead is applied to production based on direct labor-hours and its standard cost card per unit is as follows:

Direct material: 8 pounds at $10.00 per pound $80.00

Direct labor: 5 hours at $13 per hour 65.00

Variable overhead: 5 hours at $8 per hour 40.00

Total standard variable cost per unit $ 185.00

The company also established the following cost formulas for its selling expenses:

Fixed Cost per Month Variable Cost per Unit Sold

Advertising $ 290,000

Sales salaries and commissions $ 280,000 $ 21.00

Shipping expenses $ 12.00

The planning budget for March was based on producing and selling 15,000 units. However, during March the company actually produced and sold 17,000 units and incurred the following costs:

A. Purchased 170,000 pounds of raw materials at a cost of $8.00 per pound. All of this material was used in production.

B. Direct-laborers worked 64,000 hours at a rate of $14.00 per hour.

C. Total variable manufacturing overhead for the month was $513,920.

D. Total advertising, sales salaries and commissions, and shipping expenses were $300,000, $500,000, and $205,000, respectively.

Questions:

11. What is the variable overhead rate variance for March? (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance.). Input the amount as a positive value.)

12. What amounts of advertising, sales salaries and commissions, and shipping expenses would be included in the company's flexible budget for March?

13. What is the spending variance related to advertising? (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance.). Input the amount as a positive value.)

14. What is the spending variance related to sales salaries and commissions? (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance.). Input the amount as a positive value.)

15. What is the spending variance related to shipping expenses? (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance.). Input the amount as a positive value.)
Answer #1

11) Variable overhead rate variance = (standard rate-actual rate)actual hours

= (8*64000-513920)

Variable overhead rate variance = 1920 U

12) Advertising = 290,000

Sales salaries and commission = (17000*21)+280000

= 637,000

Shipping expense = 17000*112 = 204000

13) Spending variance advertising expense

= 300,000-290,000

= 10,000 U

14) Spending variance sales salaries and commission

= 637,000-500,000

= 137,000 F

15) Spending variance shipping expense

= 205,000-204,000

= 1000 F